l Verizon 7:00 PM 35% X Assignment 11-2018- T Beam Effec... Division of Engineering Technology Reinforced Concrete Design I ETC 4454 Find effective flange width b for the three beams below. 16 ft. The slabs thickness is 3 in. Assume that the span for the beams is 10 ft Beam #1 Beam 82 Beam #3 Hint: Find hy, bw and I 12 +16 befective flange widh of T- Beam beffective flange width of L-Beam eslab rhickness slab nkickness the clear dis tance to the fnce of the next webs I,:the clear dis tance to the face of the next web L - lengrh of the sup porting beam spean L-length of the sup porting beam span Open With Print
Answer #1

Beam #1:

The beam #1 is a L beam

bw=9 in

L=16 ft (Span)

hf=3 in

be=bw+L/12=9+(16x12)/12=25 in

be=bw+lw/2=9+(6x12)/12=15 in

be=bw+6hf=9+6(3)=27 in

So the effective width for beam 1 is 15 in

Beam #2

The beam 2 is a T beam

bw=8 in

be=L/4=(16x12)/4=48 in

be=bw+0.5LW1+0.5LW2=8+0.5(10x12)+0.5(6x12)=104 in

be=bw+16hf=8+16(3)=56 in

The effective width for beam 2 is 48 in

Beam #3

The beam #3 is a L beam

bw=6 in

hf=3 in

be=bw+L/12=6+(16x12)/12=22 in

be=bw+lw/2=6+(6x12)/12=12 in

be=bw+6hf=6+6(3)=26 in

So the effective width for beam 3 is 12 in